OpenStudy (russianboss1):

Fan and medal solve the system of equations x^2+y^2+2x+2y=0 x^2+y^2+4x+6y+12=0

OpenStudy (russianboss1):

@AloneS @Hayhayz

OpenStudy (tkhunny):

Have you considered simply subtracting the two equations. This gives a single line where the more complicated solutions must lie.

OpenStudy (eliesaab):

Try to draw these two circles or subtract the first equation from the second as suggested above

OpenStudy (tkhunny):

Yes, please, complete the square four times and draw the circles. THEN, draw the line given from the subtractions. You will be well on your way.

OpenStudy (russianboss1):

Is that possible, subtracting the equations? I know the solutions, I just need to know how to do them

OpenStudy (eliesaab):

If you subtract them and do some manipulations you get x=-2y-6 replace in one of the equations and use the quadratic formula

OpenStudy (eliesaab):

Replace in the first eqaution, you get \[y^2+2 y+(-2 y-6)^2+2 (-2 y-6)=0 \]

OpenStudy (tkhunny):

I'm going to guess that you have used the "Elimination Method" many times to solve small linear systems of equations. Why is this system any different? Use the same methods. x + y = 5 x - 6 = 2 Add them 2x = 7 Done.

OpenStudy (eliesaab):

Expand the equation above, you get \[ 5 y^2+22 y+24=0 \]

OpenStudy (russianboss1):

what equation did you expand?

OpenStudy (eliesaab):

Solving it you get \[ \left( \begin{array}{c} y=-\frac{12}{5} \\ y=-2 \\ \end{array} \right) \]

OpenStudy (eliesaab):

Use x=-2y-6 to find the corresponding x

OpenStudy (eliesaab):

I expanded the one in my post above

OpenStudy (eliesaab):

You will get \[ \left( \begin{array}{cc} x=-2 & y=-2 \\ x=-\frac{6}{5} & y=-\frac{12}{5} \\ \end{array} \right) \]

OpenStudy (eliesaab):

You expand \[ 2 (-6 - 2 y) + (-6 - 2 y)^2 + 2 y + y^2 \]

OpenStudy (russianboss1):

thank you !!!!!

OpenStudy (tkhunny):

What? We've never heard of an example?